OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (goformit100):

use the method of substitution

OpenStudy (anonymous):

can u show me step bby step

OpenStudy (australopithecus):

2x + y = 1 y = 1 - 2x sub that into the other equation 4x + 2y = −2 becomes 4x + 2(1 - 2x) = −2

OpenStudy (australopithecus):

now just solve for x, and then sub x into one of the equations to solve for y

OpenStudy (anonymous):

x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (australopithecus):

its just simple algebra

OpenStudy (jackellyn):

solve for y in the first equation. subtract 2x from both sides 2x-2x+y=1-2x y=1-2x subtitute into second eqn 4x+2(1-2x)=-2 distribute combine like terms solve for x and substitute again to find y.

OpenStudy (anonymous):

oooo answer is no ssolution

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

2x + y = 1 4x + 2y = −2 Using the 1st equation 2x + y=1 y=1-2x put the value of y in 2nd question, 4x+2y=-2 4x+2(1-2x)=-2 4x-4x+2=-2 you get +2=-2 so this equation has infinite solutions